Assume the following information to calculate the dollar cost of using a money market hedge to hedge 200,000 pounds of payables due in 180 days. Assume the firm has no excess cash. Assume the spot rate of the pound is $2.02, the 180-day forward rate is $2.00. The British interest rate is 5%, and the US interest rate is 4% over the 180-day period. a $388,210 O b s387.900 O c $391.210 o d s396,190 Which one is the correct answer please be sure
Answer #1
Pounds 200,000.00 A
Payable days 180 B
Spot rate for pound USD 2.02 C
180 days forward rate USD               2.00 D
British Interest rate 5% E
British Interest rate for 180 days 3% F=E/360*180
US Interest rate 4% G
US Interest rate for 180 days 2% H=G/360*180
Present Value of Pounds 195,121.95 I=A*F
Pounds to USD 394,146.34 J=I*C
Answer is $ 396,190